Q1. The new governor of California will make $4000 less than the governor of New York and $9000 more than the governor of Alaska. If the sum of these 3 salaries is $440,000 what is the salary of each governor. Solve.
Q2. In a recent year, the 3 Internet access companies with the most subscribers were Earthink, MSN, and AOL. Earthlink had 700,000 more subscribers than MSN, and AOL had 3,700,000 more than 5 times the subscribers that MSN had. If the total subscribers for these Internet access companies is 32,400,000, find the number of subscribers for each company.
Q3. The number of deaths caused by tornadoes decreased 59.2% from 1950's to the 1990's. There were 579 deaths from tornadoes in the 1990's. Find the number of deaths casued by tornadoes inthe 1950's.
Q4. If a rectangular sign is 12 inches more than twice it's height, find its dimensions if the perimeter of the sign is 312 inches.

Found 2 solutions by 303795, stanbon:
Answer by 303795(602)

Question 1
Call the California Governors salary x. It is $4000 less than the Governor of New York's salary so the Governor of New York must earn $4000 more than x (ie x + 4000). His salary is $9000 more than that of the Governor of Alaska so the Governor of Alaska earns $9000 less than x (ie x-9000).
The three salaries are x (California) x + 4000 (New York) and x-9000 (Alaska).
Add the three salaries to get x + x + 4000 + x - 9000. This totals 3x - 5000.
3x - 5000 = 440 000 (Add 5000 to each side of the equation)
3x = 445 000 (divide each side of the equation by 3)
x = 445 000 / 3
x = $148 333.33
California governor earns $148 333.33 while the New York Governor earns $4000 more ($152 333.33) and the Alaskan Governor earns $9000 less ($139 333.33)

Answer by stanbon(75887)

Q1. The new governor of California will make $4000 less than the governor of New York and $9000 more than the governor of Alaska. If the sum of these 3 salaries is $440,000 what is the salary of each governor. Solve.
-------------
Let governor of CA have a salary of "x".
Then governor of NY will have salary "x+4000"
And governor of Alaska will have salary "x-9000"
EQUATION:
x+x+4000+x-9000=440,000
3x=436000
x=$145,333.33 (salary of CA govnr)
x+4000=149,333.33 (salary of NY govnr)
x-9000=136,333.33 (salary of Alaska govnr)
--------------
Q2. In a recent year, the 3 Internet access companies with the most subscribers were Earthink, MSN, and AOL. Earthlink had 700,000 more subscribers than MSN, and AOL had 3,700,000 more than 5 times the subscribers that MSN had. If the total subscribers for these Internet access companies is 32,400,000, find the number of subscribers for each company.
Let number of MSN subscribers be "x"
Then follow the same procedure used in Question1.
--------------

Q3. The number of deaths caused by tornadoes decreased 59.2% from 1950's to the 1990's. There were 579 deaths from tornadoes in the 1990's. Find the number of deaths casued by tornadoes inthe 1950's.
---------
(# in 1950) - 59.2%(# in 1950) = (# in 1990)
(1-59.2%)(# in 1950) = 579
0.408(# in 1950) = 576
# in 1950 = 576/0.408=1411.76
# of deaths from tornadoes in 1950 = 1411.76 or 1412 people
-------------------
Q4. If a rectangular sign is 12 inches more than twice it's height, find its dimensions if the perimeter of the sign is 312 inches.
--------
Let height be "x" inches
Then length is "2x+12" inches
Perimeter= 2(length + height)=312
2(x+2x+12)=312
3x+12 = 156
3x=142
x=47 1/3 inches (width)
2x+12 = 2(47 1/3)+12= 106 2/3 inches (length)
------------
Cheers,
Stan H.